### v1

# Reviewing Intelligent Cinematography: AI research for camera-based video production

Abstract
This paper offers a comprehensive review of artificial intelligence (AI) research in the context of real camera content acquisition for entertainment purposes and is aimed at both researchers and cinematographers. Considering the breadth of computer vision research and the lack of review papers tied to intelligent cinematography (IC), this review introduces a holistic view of the IC landscape while providing the technical insight for experts across across disciplines. We preface the main discussion with technical background on generative AI, object detection, automated camera calibration and 3-D content acquisition, and link explanatory articles to assist non-technical readers. The main discussion categorizes work by four production types: General Production, Virtual Production, Live Production and Aerial Production. Note that for Virtual Production we do not discuss research relating to virtual content acquisition, including work on automated video generation, like Stable Diffusion. Within each section, we (1) sub-classify work by the technical field of research - reflected by the subsections, and (2) evaluate the trends and challenge w.r.t to each type of production. In the final chapter, we present our concluding remarks on the greater scope of IC research and outline work that we believe has significant potential to influence the whole industry. We find that work relating to virtual production has the greatest potential to impact other mediums of production, driven by the growing interest in LED volumes/stages for in-camera virtual effects (ICVFX) and automated 3-D capture for a virtual modelling of real world scenes and actors. This is the first piece of literature to offer a structured and comprehensive examination of IC research. Consequently, we address ethical and legal concerns regarding the use of creative AI involving artists, actors and the general public, in the conclusion. Learning from the implications of Deep Fake research (mainly the non-consensual use of women and children’s faces for generating pornographic content), we emphasize the importance of awareness, responsibility and accountability regarding code sharing.
keywords:
Cinematography, Vision, Graphics
1 Introduction
Intelligent cinematography (IC) research leverages artificial intelligence (AI) to assist camera-based tasks in media production. For pre production, this concerns research involved with camera planning, for example using a script [213, 70] or user input [55, 272]. For production, this includes work on automated camera-control [145, 60, 158, 26, 277], methods for content acquisition [166, 196, 296, 139] and assisting directors and camera operators [73, 146, 105, 103, 36, 27]. For post production, this involves work that avoids the need for post-production, such as through in-camera virtual effects (ICVFX) [228, 125, 76, 177], or for enhancing or diminishing camera artefacts such as light-water reflections [270]. For clarity, our definition of IC encompasses research focused on real content acquisition. This excludes research aimed at assisting post production or automated content generation using diffusion models.
IC consequently influences the worlds of cinema, gaming and televised broadcasting. The symbiotic relationship between research and video production has been historically important to the progression of video entertainment. Yet, as this relationship evolves the types of content, styles of production, and thus fields of research, have branched into separate domains. For example, we find work that automates manual capture processes [70, 139] as well as work that elevates abstract goals like enhancing audience immersion [105, 257]. This variety of objectives makes it challenging to establish a unified definition of the IC research landscape, thus limits researchers and cinematographers who wish to understand the general state and trends. Moreover, there is an absence of both general review papers and more specialised articles. Hence, in this paper, we attempt to overcome these challenges by defining and presenting a structured review on the state and future of IC research.
This review paper categorizes work by production medium. Similar to the definition of an artistic medium, this groups research by the relevant tools and practices involved with the various production types, rather than grouping work by objective. For instance, live sport and event broadcasting productions share many cinematographic practices, so are grouped under the same heading, Live Production. Doing so allows us to: (1) further sub-categorize according to related AI research fields; this provides a holistic view of the IC landscape while maintaining relevance to each production medium, (2) include research topics which, while not targeted at creative video production, have direct relevance to it, and (3) offer a view of the research landscape that cinematographers can refer and relate to. The topics covered for each medium are outlined below.
  1. 1.
    General Production: Work on human-computer interaction for camera control, automated directive assistance and workflow optimization.
  2. 2.
    Virtual Production: Work on ICVFX and virtual stage production, and automated 3-D capture of dynamic scenes and human actors.
  3. 3.
    Live Production: Work on real-time video correction, object tracking and human pose estimation for live broadcasting.
  4. 4.
    Aerial Production: Work on single UAV and mutli-UAV/swarm cinematography, covering topics such as real-time camera planning based on visual aesthetics and safety-first solutions in cluttered filming environments.
1.1 Scope and related work
The focus of this paper is camera-oriented research; hence we do not cover popular creative AI topics such as large language models (LLMs) for script generation or generative AI for video generation. We acknowledge, however, that these may become directly involved in cinematography in the future. For instance, LLMs may prove useful for text-driven camera control or workflow optimisation. Moreover, the rapid progression of generative AI models could deliver high quality video-based methods that offer an alternative solution to automated 3-D capture. As the relevant literature does not yet consider camera-based objectives tied to these technologies we do not investigate these paradigms in the main body. Instead, we provide background on generative AI and discuss future use cases of these technologies within the relevant subsections of the main body.
There are a few articles that review AI in the context of cinematography and those that do focus on special cases. For example, Christie et al [57] reviews camera control for computer graphics. This paradigm involves view point computation and motion planning systems, and considers technical constraints such as direct and assisted control, and cinematic constraints such as view composition and pose. Similarly, Chen and Carr [46] reviews autonomous methods for camera planning, control and selection for multi-camera systems. Furthermore, Qi et al [198] reviews camera calibration techniques. Other works look at use-cases for IC such as Galvane et al [80] for video editing, Young et al [281] and Sharma et al [229] for gaming and animation and Mademlis et al [156] for unmanned aerial vehicle (UAV) control. Our survey aims to bridge related works to provide a holistic view of the IC landscape so that computing and cinematographic communities mutually benefit from the discussion.
The structure of this paper is as follows. In Section 2 we introduce the technical background on AI research commonly found in IC. The main body is contained in Section 3, where the order of the subsections is reflected in the prior list. For each subsection, we present an overview of the current paradigms and approaches and conclude with a discussion on the current challenges and future works. Lastly, in Section 4, we offer our concluding remarks and outline research that we believe has the most potential moving forward. As cinematography typically involves capturing real people, we append a section on social responsibility to our conclusion to highlight the importance of ethically driven research and content capture. We continue to emphasize the importance of awareness, responsibility and accountability regarding code sharing throughout the paper.
2 Technical Background
2.1 Convolutional Neural Networks
Convolutional Neural Networks (CNNs) are a class of neural network that extract the features of an image-type by filtering through a neighbourhood of pixels with learnable filter coefficients, to learn common patterns within a set of training image(s). Regarding IC, CNNs provide backbone architectures for visual generative AI (Subsection 2.2) and object detection (Subsection 2.3).
A common application of CNNs is in image classification, where the basic structure is shown in Figure 1. First, an image is input to the first layer through $N$ channels, where each channel represents a different representation of an image. For example, an RGB image will have three values (red, green and blue) for each pixel, thus $N=3$. The outputs of this layer then feed into the hidden layers. The outputs of each layer are called feature maps. The last hidden layer transforms the feature maps to feature vectors, so that the output layer can interpret a class probability using the feature vector. An image is then classified if it’s probability falls within a predefined confidence-interval. In the case of Figure 1, if our confidence interval is $0.5$, we classify the image as containing dogs.
Figure 1: Example diagram of a CNN for image classification
The hidden layers of a CNN can be flexibly configured. The four components generally employed are convolution layer, activation function, pooling layer and fully-connected layer:
  1. 1.
    A convolution layer applies an $n\times m$ kernel (filter) over each pixel neighbourhood from an input channel. To deal with pixels on the perimeter of an image, we generally apply padding (the type of padding can vary). Dilated [284] or deformable convolutions [64] are also used to modify the structure of the kernel for special cases. In addition, a stride can be applied to skip some pixels, i.e. the filter applies to the pixels located $s$ distances from each-other ($s=1$ means all pixels are filtered and $s=2$ means every second pixel in a row and column is taken).
  2. 2.
    An activation function determines how the features of a layer are transformed into an output. To be used in a hidden layer, the activation function needs to be differentiable and nonlinear, otherwise we would not be able to calculate the gradients of parameters during a backward pass of the network. Rectified linear unit (ReLU) and LeakyReLUs layers are frequently used as they are a differentiable non-linear function, with the added benefit that they are not susceptible to vanishing gradients (where gradients of parameters become too small to make regressive change to the model’s parameters). Several types of activation function exist, as discussed in Baheti [12].
  3. 3.
    A pooling layer effectively down-samples features, retaining the more important features from a set of feature maps.
  4. 4.
    A fully-connected layer is one where each node is dependant on all the outputs from the previous layer. This allows the network to introduce a wide range of dependencies between parameters.
2.2 Generative AI
Generative AI concerns various domains, for instance text-to-video synthesis [14, 1, 143] and novel view synthesis for dynamic scenes [141, 196, 78]. As the novel view synthesis problem-landscape is broad and specially relevant, we reserve the discussion for a later subsection. In this subsection, we focus on generative AI involving 2-D images and video.
The purpose of generative AI varies drastically depending on its application. For example, it can involve learning representations for a specific medium, e.g. text or images, and translating/transforming the representation into alternative representations to produce related content e.g. images generated by prompts and vice-versa. Terms such as text-to-image or image-to-text generation, fall under the broader class referred to as sequence-to-sequence translation [254, 276, 183, 75]. Figure 2 demonstrates how abstractions such as text, images and video frames could be sequenced, and simplifies an encoder-decoder pipeline that uses pre-trained networks to inform the encoding, decoding and/or transformation process. Overall, there are four prominent architectures that contribute separately towards image and video synthesis: (1) encoder-decoder architectures, (2) generative adversarial networks (GANs) [63, 87, 123], (3) transformers [253] and (4) diffusion models [102, 291, 72]; discussed below.
Figure 2: Text, image and video frame input sequences are encoded into a feature/latent representation and translated/transformed into alternative representations that are decoded to output other tangible representations such as text, images and video frames. $s_{i}\in S_{j}$ illustrates how a sequence may be constructed from a set of input data. $\xi_{n}$ encodes the input sequence into a representation, $n$, and a function, $F$, translates the representation into $m$. $\xi_{m}^{-}1$ decodes $m$ back into a sequence. $\phi_{k}$ is a pre-trained network that can be used during the encoding, decoding and/or transformation stage.
In a GAN (see Brownlee [30]) a random vector is fed into a generator to produce a synthetic representation. Subsequently a comparison is made between the synthetic and real image, using a discriminator. The output of the discriminator indicates the level of similarity between the real and generated content, which is subsequently used in a loss function to train both the generator and discriminator. With knowledge of the vectors that successfully generate realistic content similar vectors can be constructed to produce similar content. This introduces the idea of a latent space. Consider a 4-D input vector, $v=[0.2,0.1,0.99,0.27]$ with values between 0 and 1, that produces an image of a shoe. The values, also called latent variables, are used to infer various features. For instance, the first dimension, $v_{0}=0.2$ may be responsible for modelling features relating to shoe straps and laces, where a $v_{0}=0$ returns a laced shoe, $v_{0}=1$ returns a strapped shoe and $v_{0}=0.5$ returns a shoe with a strap and a lace. Pushing the analogy further, consider the relationship between different latent variables in our input vector. For example, if $v_{1}$ represents low-top or high-top shoes, then the combination of $v_{0}$ and $v_{1}$ will produce high-tops with laces that are tied at the tongue of the shoe and high-tops with laces that are tied at the throat of the shoe. This works so long as the training data set is sufficiently diverse to learn these various features. An acknowledged limitation for this type of GAN is the correlation between the dimensionality of the input vectors (which increases with the number of features/objects to be modelled) and computation cost (including training time, hardware and energy consumed).
To overcome this problem, conditional GANs [264, 176, 236, 121] condition the random input vectors on context-specific cues, such as object classes or scene descriptors. This is accomplished by inputting the additional information into the generator and discriminator, thus facilitates the generation of highly diverse content. For example, to generate images of wild animals, the additional information could be the class of animal (e.g. frog, bear and lion).
For sequence-to-sequence translation, GANs form a part of the translation pipeline, replacing $\xi_{m}^{-1}$ in Figure 2. GANs adapt well to new sequences, although they are less robust at a local-scale; this is where transformers excel. Transformers [148, 127] are predominantly used to model the relationships between different sequences, replacing $F$ in Figure 2, and place greater attention on local context. The attention mechanism essentially equates the significance of prior values and their distance from a specific value in a sequence; this process is called self-attention. Thus, the attention and contextual value of each point in a sequence is used to generate a representation that is decoded to provide a tangible output. A familiar example is the generative pre-trained transformer (GPT) model [201, 29], popularly used for text-to-text and text-to-image translation. For example, if we have a database containing the input question “Are you ok?” and various outputs, such as “I am fine” or “I am sad” the transformer will learn the likeliest response sequence. For example on the local scale, it may be that the majority of answers begin with “I am”, so this is first selected as the start of the output. If a subsequent majority of answers follow this with “fine” and end the response here, this will make up the second part of the output. There are also cases where context drastically changes the output. For example, a similar question, “Are you actually ok?”, may illicit a meaningful reply if meaningful replies appear more frequently within the training data set. Therefore, more attention might be placed on the word “actually”, triggering a different output.
Ultimately, transformers provide a novel paradigm for text-oriented generative AI, notably machine translation [301, 108]. This is especially important to cinematographers interested in using text-controlled methods either for automated video/image analysis/editing [108, 28] or for image and video generation [160, 289]. The Stable Diffusion paradigm is closely related to this and is currently the dominant method for generating images from detailed prompts.
Stable Diffusion embodies a different approach to image and video content generation. Until recently, generative methods, such as GANs, encoder-decoders and autoencoders (described in Figure 2) have struggled to provide solutions that offer both low cost (energy, computation, etc.) and fast inference. This is the accomplishment of stable diffusion and a predominant reason for its current success. Stable Diffusion [7] works by first generating a random N-dimensional tensor - similar to the random vector generated in a GAN. The tensor, together with an encoded prompt, are input into a noise predictor (usually a U-Net [216, 231]) which outputs a new tensor predicting the noise in the input tensor. This noise tensor is subtracted from the input tensor, essentially denoising the input without reducing the resolution of the input tensor. The new tensor is re-used as an input (now without a prompt) and a new, less noisy, tensor is determined. This process repeats for a fixed number of steps, whereby the final and least noisy tensor is decoded into an image. As the step from a latent space to a less noisy space is a simple operation for a GPU to perform (i.e. a subtraction of two tensors), the generation process is fast. Additionally, with sufficient iterations, the generated image contains negligible noise and is thus of high quality.
Video-based methods using GANs [48, 144, 150], transformers [266, 227] and stable diffusion [42, 122, 41] present a number of underlying challenges, including issues with temporal consistency and a lack of suitable quality assessment metrics. Furthermore, current methods rarely consider camera-based tasks, such as the optimal camera pose or trajectory for generating a video. Considering the current pace of research, we believe it won’t be long before this is possible.
In conclusion, generative AI is a powerful tool and may offer numerous solutions to production challenges moving forward. However, with great power researchers must actively seek ethical solutions to releasing code and/or pre-trained models. While many publicly available models (or online application interfaces) incorporate safety measures into their code (like detecting malicious text-based inputs) there are still cases where this is not sufficient for avoiding pornographic content generation [214]. For researchers, steps should be taken to raise and act on concerns prior to releasing publicly available code, dataset or pre-trainedd models.
2.3 Object Detection
In many instances in IC, there’s a need not just for automatic object recognition, but also for locating these objects within an image or a video sequence. For example, in live football broadcasting, there is a need to classify player poses for the purpose of automating camera control. This can assist upcoming phases of a game or event, leading to better organisation of shots. Considering the variation of background clutter (e.g., from a crowd, grass, dynamic advertisement banners, etc.), object detection poses significant challenges. Location can be identified by either pixel-level contouring around the object’s edge or by a bounding box. The former involves pixel-level classification and segmentation, together called semantic segmentation. In contrast, while the latter incorporates a regression branch to estimate the four corners of the bounding box alongside the classification branch. Object detection using bounding boxes is faster than semantic segmentation and has been extensively utilized in real-time applications, such as sports broadcasting. For example You Only Live Once (YOLO) models for object detection and image segmentation [207] only require one pass of a network to infer location. Furthermore, image masking can help YOLO reduce stationary noise [288] by excluding known pixels regions that do not contain relevant information during inference.
YOLO works by dividing a target image/frame into $N$ cells having size $s\times s$ resolution; these are called residual blocks. Then, using bounding box regression the model estimates which residual blocks are associated with which bounding boxes. The probability of an object being present within the residual block is also predicted. Using non maximal suppression the model is able to reduce the influence of low probability scores and determine the bounding boxes with highest confidence; as illustrated in Figure 3. Following this, to measure the accuracy we take the intersection over union (IoU) of a ground truth bounding box and the estimated box to determine how close to the correct bounding box the prediction is. This provides a loss function to use during model optimisation Bandyopadhyay [15].
Figure 3: A visualisation of the stages of the YOLO process with an example of how the IoU is determined.
Several versions of YOLO have been used in IC applications, the current variant being YOLOv8 [117]. YOLOv2 supports the prediction of a fixed number objects in a single block, resulting in improved detection of smaller and grouped objects. YOLOv3 builds on this by introducing a model (with higher complexity) that can detect smaller objects and preserve fine detail. YOLOv4 [22] (proposed by different authors) adds to YOLOv3 with cross mini batch normalisation for higher accuracy and weighted residual connections for better convergence during learning.
2.4 Camera Pose Estimation
Camera pose estimation approximates the orientation, path and motion (the extrinsics) of a camera in 3-D space which, for example, is useful as a basis for mitigating motion blur caused by the motion of a camera. Accurate camera tracking also enables automated object detection, human pose detection and 3-D capture (photogrammetry). There exist numerous (commercial) solutions for this involving additional hardware (e.g. LIDAR or GPS) for local and global positioning. Though, there is also research on video-based solutions for cases where physical positioning systems are unavailable.
There are three main approaches to pose estimation: (1) Visual Odometry (VO), (2) visual Simultaneous Localisation and Mapping (vSLAM) and (3) Structure from Motion (SfM), all of which share common components [241]. The objective of VO is to recover a camera path incrementally, optimizing the current pose given the prior set of poses (window bundle adjustment) [40]. We consider VO as a short-sighted solution as it has trouble linking a poses that intersect previously visited locations, i.e. the new set of frames would be treated as a separate locations. Thus, vSLAM learns an additional global localisation map as well as inheriting VO’s objective of optimising the consistency of the local trajectory. The constraint with using vSLAM is that there are limited number of cameras and window bundle adjustments that can be processed. SfM approaches [2, 58] overcome this by utilizing measurements taken for every viewpoint/frame to boost reliability and avoid degenerative cases [194], including those in Molton and Brady [168], Mallet et al [159] and Hirschmuller et al [100]. This is a popular means of pre-processing 2D images for 3-D modelling.
Further, discussions on this topic include Saputra et al [221] who compare vSLAM and SfM methods for highly dynamic scenarios, and Taketomi et al [241] who assess vSLAM and SfM methods between 2010-2016, and Yousif et al [282] who compares state-of-the-art general VO and vSLAM methods for robotic controllers (applicable to automated camera control in IC). There are numerous use cases in IC: for example, in Section 3.4 we discuss several instances where this research is involved with aerial capture. Additionally, research on camera calibration heavily supports work on automated 3-D/volumetric capture, as the ability to position images in 3-D space is heavily tied to the quality of inverse rendering pipelines [147, 53, 16]. This discussed further in the following subsection and is revisited in the main body.
2.5 Automated 3-D Capture
Automated 3-D capture is a popular topic in IC research, whether for analysis or for solving the inverse graphics problem. Here, the objective is to discover a 3-D representation using 2-D imagery and/or other sensor data. In Figure 4, we illustrate simplified representations of the current approaches to the problem.
(a) Photogrammetry relies on sensor data (may include cameras) to estimate a point cloud representation of a scene
(b) Neural Radiance Fields sample points in space to estimate a colour and density value for each sample projected along a pixel-ray
(c) Gaussian Splatting uses a point cloud to estimate the position of samples in space and uses this to estimate the colour, density scale and rotation of each Gaussian point
Figure 4: Simplified representations of current approaches to automated 3-D capture
A popular capture method is Laser Informed Photogrammetry (e.g. LiDAR) which creates 3-D point-clouds from measurements of point depth relative to a camera’s position. This can now be done on a mobile phone thanks to the introduction of mobile LiDAR camera sensors. These methods are promising for capturing enclosed spaces, however there are challenges with processing information over greater distances as well as limitations associated with noisy data. Additionally, these cases require photogammetric sensors which limits applicability when only image/video data is available.
Recently, neural radiance field (NeRF) methods have been used to solve this problem, whereby the aim is to model the rendering processes involved with automated capture. Unlike Photogammtery which may represent scenes as 3-D structures (mesh, voxel-grid, point cloud), neural methods represent scene properties as neural representations which are sampled to interpolate views defined by the camera extrinsics in a virtual space. Given a set of images describing a real environment, NeRF networks are expected to reliably learn the visual features of a scene.
NeRFs represent an important breakthrough for cinematographers, as can be adapted to a range of cinematographic tasks. For example, rendering 3-D from 2-D images can avoid costly re-shoots caused by poor lighting, scenery, weather and acting deficiencies [167, 255, 287]. Shots can also be re-worked in post production, meaning that shot-type, camera jitter, pacing and focus can also be modified. While the field of NeRF research is relatively new, it has developed rapidly and can now deliver high quality, compact solutions, leading us to expect production-ready NeRFs in the near future.
The NeRF model samples volume density, $\sigma$, and colour radiance, $\mathbf{c}$, provided a 5-D input comprised of 3-D co-ordinates, $\mathbf{o}\;\epsilon\;\mathbb{R}^{3}$, plus 2-D viewing direction, $\mathbf{d}\;\epsilon\;\mathbb{R}^{2}$, which represents the position and viewing direction of a sample in space. A sample can be thought of as either a volumetric line-segment along a ray or a voxel intersected by a ray casted into the scene. Simply put, for each pixel $(x,y)$ in an image a ray vector $\mathbf{r_{x,y}}=\mathbf{o}+t\mathbf{d_{x,y}}$ exists, where $t=0$ represents the focal point of an image, $t=n$ represents the position of the image plane along the ray and $t>n$ represents point samples along a ray where $n$ sets the scalar distance from the focal point to the image plane and can vary w.r.t lens distortion. We refer the reader to the Nerfstudio documentation [243] [^1] Accessible online: https://docs.nerf.studio/en/latest/nerfology/model_components/visualize_cameras.html which overviews the different types of camera models, sampling schemes and sample representations that are found in the NeRF literature.
The original NeRF paper [166] defines the network as a multi-layered perceptron (MLP) with inputs $(r_{i},d_{i})$ for each sample in a bounded space $t_{near}<t_{j}<t_{far}$ and outputs $(c_{i},\sigma_{i})$. To render samples aggregated along a given ray, Mildenhall et al [166] proposes Equation 1 where the exponent represents the accumulated transmittance w.r.t to the volumetric density of preceding samples. In practice, Equation 1 is numerically approximated using quadrature, in Equation 2, where $\delta_{i}$ is the thickness of a volume sample along a line segment.
[…]
$\hat{\mathbf{C}}(\mathbf{r})=\sum^{t_{far}}_{i={t_{near}}}(1-\exp(-\sigma_{i}\delta_{i}))\mathbf{c_{i}}\exp({\sum^{i-1}_{j=t_{near}}-\sigma_{j}\delta_{j}})$ (2)
Subsequently, a loss function $\mathcal{L}(\mathbf{C^{*}},\hat{\mathbf{C}})$ is used to optimise the predicted ray colour w.r.t to the colour of a ground truth pixel, $\mathbf{C^{*}}$.
To reduce the influence from spectral bias [202] NeRF maps position and viewing directions to a view-point’s fourier features, using the encoding $\gamma$ in Equation 3, where $k$ is a hyper-parameter defining the dimensionality of the feature vector (the bandwidth of our discretized frequency encoding). It is well known in neural representation research that coordinate-MLPs struggle to learn high-frequency signal details, hence the need for encoding frequencies using $\gamma$. However there are several studies that discuss alternative changes to the MLP activation functions which forgo the need for discretized frequency encoding [233, 224]. For example, Saragadam et al [224] proposes an MLP using wavelet activation (called the wavelet implicit representation representation (WIRE)) while Sitzmann et al [233] proposes a sinusoidal activation (called sinusoidal implicit representation (SIREN)). These are shown to not only reduce the size of the MLP but also capture a higher-bandwidth of frequencies. However, despite Sitzmann et al [233] and Saragadam et al [224] showcasing higher quality and faster convergence, they discuss how un-actuated these results are in NeRF research.
$\gamma^{k}:\mathbf{p}\rightarrow(sin(2^{0}\mathbf{p}),cos(2^{0}\mathbf{p}),\cdots,sin(2^{k}\mathbf{p}),cos(2^{k}\mathbf{p}))$ (3)
Several approaches have been proposed to speed up NeRF computation: multi-resolution hash encodings, voxel grids and voxel-trees [171, 262, 283, 133]. This is particularly relevant for researchers interested in the dynamic NeRF paradigm, as model complexity increases significantly when modelling 4-D space and time. The PlenOctrees representation [283], illustrated in Figure 5, builds off NeRF’s ability to functionally represent a view-dependant scene by, (1) representing the structure of a scene as a sparse-voxel octree [268, 220, 133], and (2) sampling colour by involving spherical harmonics. (1) facilitates the rendering step through fast-to-access representations (octrees). Whereas (2) allows us to model view-dependency as octree leaves by mapping the view dependant colour values to a sphere’s surface and predicting the coefficients of the spherical harmonic equation (with fixed degree) to return a colour value where the pixel-ray intersects the sphere. Subsequently, the rendering speed and learning is optimised significantly, and one could argue novel-view quality has improved too. One caveat is the amount of data a representation will consume (in the magnitude of gigabytes), which perhaps is acceptable for those with required hardware, however there exist more affordable explicit options [45, 212]. Additionally, there is some difficulty with visualising unbounded scenes as voxel-grids occupy finite space. Yu et al [283] overcomes this by using NeRF++ [290] for rendering out-of-bound scenery. This is accomplished by modelling the foreground and background as separate components.
Figure 5: Method for neural rendering of 3-D images presented in Yu et al [283]
With similar motivations, 3-D Gaussian Splatting (GS) has been proposed by Kerbl et al [126] as a means of significantly reducing rendering time by using: (1) a point cloud representation of Gaussian “blobs” with position, covariance, colour and opacity properties, reducing the unnecessary computation of empty space present in NeRF models and (2) tile splatting for rendering. The covariance property is represented as a $3\times 3$ matrix $\Sigma$ in Equation 4 and determines the scale $S$ and rotation $R$ of the blob in space. In practice, $\Sigma$ is approximated to a $2\times 2$ matrix, avoiding the last row and column in the original matrix, leading to faster computation.
$\Sigma=RSS^{T}R^{T}$ (4)
Additionally tile splatting also speeds up rendering with the following steps. We highlight the importance of the sorting approach, (3), as this limits NeRFs from achieving similar computational goals.
  1. 1.
    Divide a rendering view into $16\times 16$ tile
  2. 2.
    Cull blobs with $<1\%$ confidence of intersecting each tile frustum or blobs that fall outside the near and far bounds of the camera
  3. 3.
    Sort blob depth w.r.t to each tile using a fast GPU Radix sort [164]; not per pixel as done in NeRF.
  4. 4.
    Render pixels w.r.t the sorted blobs for each tile by $\alpha$-blending until the accumulated $\alpha$ for each ray becomes 1
Finally, to train a GS model a scene is initialised with a sparse set of point-clouds that are “densified” during training by cloning, re-scaling and re-positioning blobs to fit the geometry corresponding to a set of training view/image. Blobs that are essentially transparent (low $\alpha$ contribution) are pruned.
Conclusively, GS maybe selected over NeRFs due to their enhanced computational abilities, and in many cases improved performance. However for the dynamic paradigm, the current set of benchmark scenes only last up to a minute. As existing dynamic GS models are much less compact than NeRF alternatives, this presents a challenge for storing and sharing 4-D models. Additionally, as GS is a more recent development than NeRF we are yet to see the same attention towards cinematographic based tools, such as scene and camera editing or representational transformations such as GS to mesh. Though, the research space is adapting at a fast pace so we believe it is only a matter of time before we see these limitations addressed.
Finally, we underpin the importance of social awareness and responsibility involved with this field. Unlike generative AI, we notice almost no attention to the ethical problems that may arise from creating digital doubles of real humans. We urge researchers involved with this field to take more of an active role raising awareness.
3 Intelligent Cinematography in Production
3.1 General Production Applications
In this section we present research relating to camera management and control, visual analysis, cinematographic assistance and workflow optimization.
3.1.1 Computational Language Structures for Scene Analysis, Automated Labelling Schemes and Camera Management
Cinematographic language is used to communicate the state of production and relevant processes. It enables production staff to effectively communicate and react coherently to unforeseen dilemma. IC applications require a similar level of attention. Cinematic theory is often formalized for tasks such as automated camera control and scene analysis. For example, in Jhala and Young [115] state machines are used to model cinematographic shots for camera planning in a virtual environment. In this context, in this section we discuss current approaches to human-AI interaction, with particular reference to automated camera control; we refer to this as idiomatic capture.
A heuristic approach to camera control was introduced by Jhala and Young [114], based on the decompositional partial order causal link planning algorithm [280]. This formalizes idiomatic capture by linking an approximated scene representation to a predetermined set of camera control responses. A conflict resolution algorithm ranks known actions with conditional operators to determine the optimal action and respective duration. To express idioms, Jhala and Young [114] distinguishes a set of four requirements to mediate contextual differences between different productions: (1) Story Representation, (2) Real World Representation, (3) Rhetorical Coherence and (4) Temporal Consistency. The first two points describe the physical and contextual landscapes of a set, hence it is necessary to have both geometric (or visual) and semantic representations of a production. The third point expresses the need for rhetorical structure to ensure actions are executed decisively, e.g. a hierarchical-model for selecting shot-types. The final point specifies consistency with regards to the temporal aspect of filming.
Similarly, the Declarative Camera Control Language (DCCL) uses a heuristic-based decision tree for idiomatic capture [55], Figure 6. This method sets forth a hierarchical structure for automated shot-composition and camera control by breaking down scenes into idiom-specific frame sequences. Here, the relationship of consecutive frame sequences is dependant on temporal links between idioms. A heuristic evaluator then selects a candidate action by scoring possible responses and evaluating decisions via a decision tree.
Figure 6: A three stage pipeline which uses idioms defined by camera and actor states as inputs to a DCCL compiler. Selected frames are evaluated w.r.t to the “film tree”
Idiomatic language structures have also been used for shot labelling [215] and evaluation [285]. In more recent times, the Prose Storyboard Language (PSL) [215] has been proposed as a high-level language for describing shots relative to visual composition and idiomatic camera control. This is demonstrated with Alfred Hitchcock’s North by Northwest, where a prompt not only describes the shot-type and fore, mid and background compositions (separately) but the type of transition as well. As PSL uses AND/OR decision trees it can be easily configured to add/modify labels. While the authors have not tested this on an automated labelling scheme, there is cause for such investigation as idiom-based languages might overlook newly adopted idioms and foreign capture practices.
Embedded Constrained Patterns (ECP) is a dedicated query language, conscious of physical cinematographic limitations [272]. Unlike PSL, ECP is more comprehensive in its description - labelling montages rather than individual shots. Labels are assigned using the following descriptors:
  1. 1.
    Framing Constraints: size of actors and objects, angle of shot, region of actor and object locations and camera movement
  2. 2.
    Shot Relation: size, angle and region relative to a sequence of shots
  3. 3.
    Sub sequences: a local grouping of related shot sequences
Wu and Christie [272] propose ECP alongside a automated search algorithm for optimal shot-sequence description. The search method is partitioned into two stages, (1) build a cache of valid solutions, and (2) apply a double recursive depth search to choose optimal description from the set of valid solutions. During stage (1) valid solutions are separated into three sets: (i) FC - a set of frames which satisfies framing constraints defined by ECP, (ii) RC - a set of frame couples, $[f_{i},f_{i+1}]$ which satisfies the ECP’s relational constraint, and (iii) SC - a set of frame sequences, $s_{m}=[f_{l},f_{F}]$, where $f_{l}$ and $f_{F}$ represent the start and end frames of the frame sequence. In the second stage, the search process iterates over all frames in each sequence, $f_{i}\epsilon s$, where subsequent frames are validated as part of the local sequence, $s$, or global sequence, $S$.
Alongside the technical implementation of ECP, the authors have plans for integration with Unity 3-D to apply ECP to their montages; which we see as a positive yet infrequent consideration in IC. It would be interesting to reverse the order of use. Whereby, rather than providing linguistic descriptors of a 3-D environment we could trigger camera actions through prompts written in PSL or ECP.
Additionally, we find that languages built for human-AI interaction are not easily comparable. The choice of language structure can vary drastically depending on the task for automation. While there is no current standard for general cinematography, we observe interest from Movie Labs, who look to define a set of computational linguistic structures as a standard for cloud-based workflow optimization [132]. The shift to cloud-based computing is indicative of another paradigm of production. Notably, as assets and tools are usually stored or executed offline changes to shared resources can be difficult to monitor on one platform. This could be facilitated by file management approaches which mimic, for example, code-sharing practices. However, this may be difficult to scale for productions requiring a large number of assets or productions which involve external international collaborators who may be un-knowledgeable of standard practices.
3.1.2 Directive Assistants
Converting most cinematographic concepts into controllers presents a challenging task [56]. Notions of shot compositions, shot types and shot transitions are bound by real world problems such as the cost of production, physics and the topology of a set. Thus, AI directive assistants (DAs) have been introduced to alleviate these production challenges. Current IC research addresses tasks such as deriving shot lists, shot plans, optimising camera placement and controlling robotic camera rigs.
There is a tendency to using semantic representations of a work environment to provoke DAs. For example, through DCCL, Christianson et al [55] demonstrates a heuristic approach that relies on idiomatic-based practice for decision making on shot composition. Similarly, He et al [94] discuss a different heuristic approach, utilizing a set of finite state machines [^2] A state-based control architecture for decision making. If approached intelligently (e.g. state labels represented by a bit-string), logical optimization can be applied through state transition tables and state maps/implication charts. More on this here: https://inst.eecs.berkeley.edu/~cs150/sp00/classnotes/katz-ch9-mod.pdf. to handle idiomatic camera actions. Both of these methods use virtual simulations with user-led events to demonstrate the ability of their tools to derive an idiomatic responses in a short amount of time.
There are more novel approaches to the problem of deriving a shot-list, such as de Lima et al [146] who employ an architecture that simulates four critical filming roles: (1) a Script Writer who observes the context of a current scene and sends information to (2), (2) a Screen-orgrapher who configures the staging of actors and objects for dramatic effect and passes this information to (3) and (4), (3) a Director who extracts important information and uses multiple support vector machines (SVMs) to make decisions on selected shots, and (4) a Cameraman who follows idioms for shooting. This method compounds three 2-D feature matrices into a high dimensional feature space - selecting the most relevant SVM from a pre-trained set to classify the optimal shot position and viewing angle. SVMs work by using kernel maps to remap all features into a high-dimensional matrix. This is done under the assumption that classes of features are not linearly separable until they are compounded into a high-dimension. In practice, normalized positional values represent the environment features, while the perceived emotional state of each actor is represented by the actors’ features, and the actor who is the principal focus is represented within the scene’s features. These are used to for selecting the right SVM to estimate the optimal camera location and angle given the context of the scene. As SVMs may be considered old w.r.t to the current literature, there are likely more suitable approaches for modelling the environment, actor and scene features to classify the optimal camera angle and location from a pre-defined set. For example, with a large enough data set a deep neural networks will generalize better. There are also faster and simpler approaches such as using K-nearest neighbours on the environment, scene and actor features.
Alternatively, CamDroid is a well known state-based camera control architecture, that was introduced by Drucker and Zeltzer [73] as illustrated in Figure 7. This method uses Tool Command Language (Tcl)  [^3] A high-level functional programming language. with TK embeddable language widgets [180] to interface between user-input and controllers, and pre-specified scripts. The pre-specified scripts access an object database (containing information about the local real environment) through application specific processes/object interfaces, and camera modules through camera interfaces; which ultimately determine the camera state and its subsequent action - dependant on a set of filming constraints such as the type of camera motion. Conditional functional frameworks like CamDroid are a classical way of handling automated camera control and are powerful for capturing real-time time-dependant action.
Figure 7: A diagram of the CamDroid architecture, Drucker and Zeltzer [73]
Despite the sparsity of DA research, there are recent works such as Stoll et al [240] who investigate shot selection with a multi-view camera setting for filming theatrical performances. The authors record performances from multiple views in 4K and by cropping the high-resolution frames into lower resolution frames, a set of camera actions is derived. Subsequently, skeletal and facial poses are estimated for each actor in the set of cropped videos. These are used in an automated editing script where, for example, moving lips may be used to detect an actor as the principle focus of the scene, so the relevant clip (i.e. a selected view and camera action) are selected for a given time frame. Preceding Stoll et al [240], prior works have tackled zoom and crop methods in different settings [81, 82, 199] such as Kumar et al [131] who displays the set of cropped shots in a split screen for shot selection.
The principal areas for investigating DAs: through idiomatic language, classification and state based approaches, are not novel by today’s standards of AI research, but are nonetheless capable of achieving reliable idiomatic camera actions in response to user input. However, because the discussion on concurrent methods is limited by the sparsity of recent work, we have yet to see promising work that moves away from offline automated DA and towards a live setting, such as in sport broadcasting. Despite this, numerous related fields approach similar paradigms. For instance Besançon et al [20] discuss state-of-the-art approaches for interacting with 3-D visualizations. The authors outline numerous areas for improvement, particularly surrounding the tools and error metrics used to understand and evaluate human-computer interaction. Skeletal estimating and facial recognition, such as in Stoll et al [240], are not sufficient to gain the context of a scene. Likewise, Ali [4] discuss applications relating to camera control in virtual environments, which also relates to the discussions in Sections 3.2.2 and 3.2.3.
In the following subsection we present recent approaches to workflow optimization that encompasses aspects of the DA paradigm, such as language-driven shot selection.
3.1.3 Workflow Optimization and Automated Shot Composition
Pre-producton and production workflows will vary drastically as a result of budget constraints, delivery deadlines and creative objectives. Despite this, there are aspects to film-making which remain constant, including set design, the object and background staging processes, and film capture and editing. In industry, notable efforts from Movie Labs (discussed in Section 3.1.1), illustrates the potential of a cloud-based platform for streamlining production, entailing new computational language structures, methods for cloud-security and collaboration workflows. Movie Labs have produced a set of white papers [132] detailing their ambitions over coming next decade.
Machinima production (MP) for IC leverages intelligent script writing to generate camera poses for shots (tested on a virtual scene) [^4] Further discussions: Riedl et al [213] and Elson and Riedl [74].. Rather than automating the entire production process, we see potential in adapting MP tools to attend to the general scope of workflows (not necessarily for computer generated animation). For example, the GLAMOUR system [70] utilizes natural language generation informed by cinematographic idioms, to produce a movie-like composition of shots, with comprehensive descriptors of the scene from still images. The outcome being several short documentary-style productions. GLAMOUR is a multi-objective heuristic approach for attention optimization. Anantrasirichai and Bull [6], thus as animated actors are directed (by an AI) during a scene, the optimal choice of camera and transition are determined. As discussed by Tan [242], the level of potential stimuli one can experience from a film far surpasses the complexity of audience attention. Thus, GLAMOUR could be extended to include other processes involved in automated shot composition, such as kernalized filter control (KCF), presented by Henriques et al [99] and used for actor detection and framing in automated aerial content acquisition (discussed in Section 3.4.1). Additionally useful, cinematic motion style transfer in Wang et al [265] uses 3-D representations generated from real images and matched to a shot sequence from a scene in a given movie. This could be paired with attention optimization to improve upon selected idiomatic shot to transfer, from which we could derive more meaningful camera poses and motions.
A semi-automated method introduced by Yu et al [285] was designed to handle object staging, automated camera positioning and shot composition, from an annotated script. The framework shown in Figure 8 optimizes camera parameters dependant on a generated sequence of actions for each present character. An action list is denoted by $\{a_{i}|i=1,2,...,N\}$, where $a_{i}$ is the $i$th action and $N$ is the total number of actions available. Following this, the action list is transformed to stage performance, where each action corresponds to a movement (time period) in a scene. This automates scene scheduling process. The consequent performance of an action within a scene is denoted by $\{p_{(t)_{a_{i}}},p_{(t+1)_{a_{i}}},...,p_{(t+l)_{a_{i}}}\}$, where $t$ denotes the moment the movement begins for $t\epsilon T$ and $(t+l)_{a_{i}}$, which denotes the sub-components of a movement for an action $a_{i}$ within a performance, $\{p_{t}|p_{1},p_{2},...,p_{T}\}$. During the camera optimization step, “aestethic” and “fidelity” models are jointly applied to a performance $p_{i}$, to determine the optimal camera $c_{t}$ to use at each time step $t$. The aesthetic model analyses six factors for camera planning: character visibility, character action, camera configuration, screen continuity (relative to character position), moving continuity (accounting for on-going changes between movements) and shot duration. The fidelity model first assumes that a mathematical model can approximate the relationship between a script and a generated video. To accomplish this, the model uses the global vector for word (GloVe) embedding model [186] to generate text from the generated video, then analyses the similarities between generated script and target script [^5] The final result is visualized here: https://www.youtube.com/watch?v=0PUdV6OeMac.. A GloVe embedding is vector representation for words that are trained using the global co-occurrence word-word of word pairs/groups. Thus, sub-structures of the vector space can define synonyms (proximal parallel vectors) and canonical structures (vector paths). This is similar to the latent variables for text-based generative AI, discussed in Section 2.2.
Figure 8: Illustration of the pipeline for text2animation in Yu et al [285], where $c_{j}$ is the camera selected for capturing action $a_{i}$.
Overall, both of these methods produce naive approaches to workflow optimization which is constrained by cinematographic understanding and current technological capability. For example, the crux of the evaluation mechanism in Yu et al [285] depends not only on the assumption that a fidelity model can be constructed, but also the ability to re-translate a video into text in a manner appropriate for evaluation. The reliance on this chained processes reduces the effectiveness of this model. Though as Yu et al [285] concludes, there is still much to accomplish in optimising production workflows. For example, a further extension of method like GLAMOUR could involve automating the generation of annotated shots or for extending story-board process. This could prove beneficial for productions with tighter budgets and/or deadlines. Additionally, an improved method of evaluating such models is necessary, as we have doubts as to the validity of text-based testing for cinematographic models.
Considering that workflows are often fragmented due to reliance on external collaborators, research focused on supporting collaboration seems more plausible. For example, a model could be tasked with acting as an administrator for a shared resource - deciding if committed changes should be accepted given some camera-based constrains. Leaning into ICVFX, cameras such as the Axibo PT4666Product information: https://www.axibo.com/product/pt4. motion controlled slider host a cloud-based Unreal Engine workflow and directly attend to the shared resource paradigm. Use of such tools could be extended to investigations projecting dynamic assets [65, 134, 116] and validating the in-camera (artistic) composition given a feature-based target [138, 265].
3.1.4 Challenges and Future Work
One of the challenges with introducing novel AI to the production process is testing and validation. Unfortunately, real production environments are not only challenging to access but difficult to control. Therefore most proposed research favours testing in simulation through virtual environments. Clearly, there is some disparity between a semi-informed simulation and a real use case. Going forward, we need to acknowledge the copious benefit of experimental productions [91]. Experimental productions could become a platform for testing IC tools, whether this is to accomplish menial tasks so artists can focus on creative experimentation or to explicitly support one’s creative expression. Alternatively, we could envision using 3-D reconstructions of real scenes, discussed in Section 3.2, to achieve more reliable tests on real scenarios. This would provide more relatable results for cinematographers that intend using IC to direct cameras in real environments, as well as those that wish capture cinematic footage using 3-D reconstructions.
With respect to supporting the cinematographic language, there are many considerations to make. For example, the application of language can be beneficial for semantic script analysis/generation [101, 71, 161] but it could may also act as a form of communication between production staff and AI interfaces [55, 285]. We found PSL to be particularly interesting as it is represented as mid-level language. This is beneficial as it reduces the reliance of solving natural language processing (NLP) problems (e.g. a sequence2sequence encoding/decoding problem), which can be troublesome when confronted with complex shot descriptions or tasked with describing subjective cinematographic observations. Though, we should acknowledge the progress made in NLP and semantic analysis research over the last decade [34, 44]. Consequently, we are interested to see more proposals for learning linguistic structures for cinematographic production. For example, riding off recent breakthroughs in NLP [200, 165] researchers could explore LLMs for camera control and workflow optimization tasks. Consequently, we believe the progression of this will rely on establishing reliable sources of data. For example, for semantic shot analysis one could propose a comprehensive data set and method of labelling shots using PSL to reduce the complexity of natural language structures for word-embeddings.
Otherwise, DA research provides a platform to host a range of AI controllers. They can be adapted for specific purpose and architectures are often implemented as modular pipelines. However, evaluating performance in-the-wild is limited by access to real sets, production staff and the limited number of benchmark datasets and models. This indicates that the state of DA research still has a considerable journey before solutions can be applied on a commercial scale. We hypothesize that as production is further facilitated in other areas, such as improved camera control [5, 57, 114], DA could be facilitated by controlling a set of automated tasks on a higher-plane of abstraction, perhaps through the use of semantic query languages for resource descriptive frameworks (RDF) [62, 89] or DCCL and PSL.
With respect to workflow optimization strategies, we primarily question the social benefit of MP research. As discussed in Section 1, IC relies on easing the production of artistic film-making, so automating the entire process (‘text-to animation’) leaves little room for creative input [269, 192, 8]. However, this could be used to support the story-boarding process and other pre-production tasks. Taking this further, one could formulate a storyboard as a set of animated clips to make decisions on shot composition. Thus future work could look at optimising shot composition, such as for style transfer [265] or attention optimization [252, 188].
3.2 Virtual Production
In this section we focus on virtual production using real cameras. This is important as it delineates from classical virtual production used in animation or gaming, whereby two notable differences exist. Firstly, compositing virtual scenes for animation is less challenging than compositing virtual assets in a real scene for an IC production as nuances in lighting, colouring and perspective between real and virtual assets require attention. Secondly, real cameras are limited by physical and fiscal constraints, like set topology, additional hardware and technical skill for achieving specific camera motions. These matters are trivial to accomplish with a virtual camera that has 6DoF.
These differences underpin the general concerns of research tied to virtual production for IC. Hence, in this section we look at research involved with ICVFX and LED Volumes - touching upon works that investigate re-colourization and image-based lighting (IBL). We also discuss research focused on synthesizing virtual replicas of real actors and scenes through NeRFs, prefaced in Section 2.5. This looks at removing the physical and fiscal limitations of using real cameras for content acquisition, allowing users to re-capture real scenes in the context of a virtual environment/engine, as is achieved with classical virtual production.
3.2.1 ICVFX and LED Volumes
Numerous scholarly works have undertaken the task of dissecting visual effects (VFX) and in-camera applications within the realm of cinematography. However, the relevant use case(s) for IC is vaguely defined. Considering a purely cinematographic perspective, ICVFX usually involves CGI and/or compositing techniques that are executed in real time, providing cinematographers with a live feed of how virtual effects will appear relative to the real scene set-up. This can support production in numerous ways, such as an indicator for poor lighting or as a way to pre-visualize compositing and VFX to ease post production challenges. Here, the technical practice mainly concerns the ability to distinguish foreground and background elements [228] as well as modify lighting and colour [18]. Sharma [228] briefly presents chroma key and roto scoping as paradigms, with a focus on the former; in Figure 10(a). While, Bengtsson and Kang [18] presents how IBL, a well understood method in VFX in Figure 9, has evolved into a lighting workflow surrounding LED video screens for driving in-camera relighting and colourization, Figure 10(b). This hints at two distinct use cases for IC research:
  1. 1.
    Automatic foreground segmentation [^7] Here, the terms segmentation and roto scoping can be used interchangeably.: (i) with and (ii) without chroma screens (i.e. blue/green screens)
  2. 2.
    Automatic IBL and re-colourization
Figure 9: IBL maps an environment image onto a 3-D primitive (e.g. a sphere). The IBL environment is treated as a light source and using ray tracing a model is composited.
(a) Foreground separation with automated segmentation or chroma screens.
(b) In-camera compositing with an LED stage
Figure 10: Illustrating ICVFX compositing approaches
It is generally known that use case (1) is readily accomplished by using chroma screens (1.i) and keying out the colour corresponding to the screen. However this introduces issues for colour-based image segmentation. Cheng et al [50] touches upon relevant colour-based image segmentation techniques angled towards the wider AI audience. The authors discuss solutions such as histogram thresholding [149], using binary trees to store data intensive 3-D colour spaces [225, 222], region-based methods [249, 49], fuzzy techniques [110, 247] and neural networks [106, 35]. Consequently the authors highlight problems with shading (e.g. shadows and highlights) and texturing. The most apparent case of this is interfering light-bounce from the chroma screen. This leads to subject specific approaches, for example Sigal et al [232] approaches real-time skin segmentation for video opposed by time-varying illumination. Sigal et al [232] optimizes a second order Markov model to predict a skin-colour histogram over time. Additionally, chroma keying restricts the ability to present colours similar to that of the chroma screen.
More recently, we see approaches attending to case (1.ii). This is a popular paradigm outside of cinematography whereby we see emerging research such as the Segment Anything Model (SAM) [129] [^8] Still in pre-print. - a one-click solution to general image segmentation. Angled towards cinematography we find work such as Roto++ [142] - a rotoscoping tool with the ambition of respecting the artists’ requirements. Roto++ improves upon traditional interpolation techniques by combining a real-time appearance tracking and a novel shape manifold regularization process (built on the Gaussian process latent variable model (GP-LVM) [135]). Subsequently, within a sequence of frames the method (1) predicts the change of a shape manifold and (2) identifies which next keyframe needs to be manually roto scoped.
Concerning case (2); while IBL is classically a 3-D rendering technique, it can be applied to real scenes where the lighting set-up can be readily changed [67]. Ren et al [209] and Wang et al [259] discuss approaches that reconstruct the light transport matrix (LTM) [^9] This defines light interactions on object surfaces.. Wang et al [259] classifies the approaches into three categories: (i) brute force [179, 92], directly modelling the LTM, (ii) sparsity based [84, 162, 205], modelling a set of basis functions under the assumption that each row of the LTM can be linearly approximated, and (iii) coherence based [181, 259, 79], analysing the coherence reflectance field to acquire the LTM. The limitation with these methods is that they require multiple images under varying lighting condition, meaning for video this problem is more challenging to address.
Interestingly focus on cases (1) and (2) has shifted towards using LED panels to project virtual backgrounds as a practical solution. For case (1) foreground-background separation is made easy by replacing chroma screens with virtual backgrounds displayed live on interconnected LED panels. While more costly and energy intensive, it reduces the workload for post-production and avoids the need to roto scope. Unfortunately this means the need for research in this area is minimal and shifts toward supporting computation (e.g. reducing energy consumption). For case (2) the outcome is not so severe. Instead, the introduction of LED panels offers new possibilities for automated lighting calibration, now including the LED panels as a light source [174, 185, 112, 96]. For example, LeGendre et al [139] treats the panel lighting as ambient light, producing an example result in Figure 11
Figure 11: Results from images taken in natural lighting (left) and configuring the LED panel for lighting (right) provided by LeGendre et al [139]
[139] accomplishes this by first applying two matrices, $M$ and $N=MQ^{-1}$ to the out-of-camera-frustum and in-camera-frustum, respectively, where $M$ and $N$ represents a 3$\times$3 pre-correction matrix. Then applying the post-correction matrix, $Q$, to the final image, where $Q$ represents a 3$\times$3 post-correction matrix that re-maps viewed pixels to the desired/expected colour schema. $M$ is solved through matrix calculation, from known LED emission spectral sensitivity functions, i.e. $M=[SL]^{-1}$, where $[SL]$ represents the observed average pixel values from capturing light emitted by the LED panels. $Q$ is found by minimising the squared error between predicted pixel values and target pixel values, using the 3x3 matrix $[SRL]_{j}$. This encodes the spectral modulation and integrates the camera spectral sensitivity functions and LED emission and material reflectance spectra [139] for a given colour chart [^10] A colour chart is used to map shades and tints of red, green and blue, where each square in a chart is a different shade. square, $j$. Testing on colour charts showed near-optimal results, though resulting errors are limited by using a 3$\times$3 linear transformation kernel. LeGendre et al [139] highlights issues resulting in de-saturation of skin colour [^11] The authors propose further testing on a larger spectrum of skin colours. and fabrics; which is reasoned by the restrictive colour schema from LED panel lighting.
Similarly, Smith and Zink [234] looks at correcting dynamic in-camera-frustum hue changes, though using simpler colour transforms. Additionally, Debevec and LeGendre [68] investigates a method for HDR-Image (HDRI) lighting reproduction, going from virtual HDRI setting to a LED volume setting. This essentially inverts the classical IBL problem and is approached by dilating pixels above a given threshold to meet constrains on local average pixel values displayed displayed on a virtual LED wall.
Overall, the use of LED panelling in production is still new. Aside from discussing lighting and colourization, the IC landscape outside of this paradigm is undefined [125]. Subsequently, we have aggregated a set of non-scientific resources, in Table 1, which contribute to cinematographic discussions concerning the use of LED panels. We additionally categorized the list to provide further clarity on subject materials.
Table 1: Supplementary Online Resources containing articles, documentation and video-based discussions.
Topic Papers
General Discussions Nila et al [174], Kavakli and Cremona [125], Kadner [120], Pires et al [189], Kadner [119], Hendricks [97], Society [235]
Workflows/Pipelines Chambers et al [43], Consulting [61]
Motion Tracking Viehmann [256], Televisual [244]
Actor Imersion Bennett [19], Leane [137]
3.2.2 Camera Calibration and Localization
The objective of camera calibration and localization (or pose estimation) is to map a 3-D world onto a 2-D image plane. This involves modelling camera intrinsic/inertial parameters and extrinsic/external parameters [295] using real image data. An illustration of the relevant parameters to be modelled is shown in Figure 12. The intrinsic parameters define the cameras model, focal length, and lens distortions. The extrinsic parameters model the camera transform matrix for each pose as well as the path for moving shots.
Figure 12: Extrinsic parameters are shown in grey. Intrinsic parameters are shown in white.
This is a popular problem that concerns the general use case of a single camera [198, 152, 208] as well as specific use cases, such as surgical monitoring [198, 175, 130]. Regarding IC, we find work supporting a number of applications including aerial photography [25, 195], photogrammetry [154, 59, 77], image-based 3-D reconstruction [47, 250, 275, 170] and underwater filming [230, 38, 155, 163, 299]. The two practical uses for this in cinematography are automated camera control and real 3-D reconstruction.
For automated camera control, the research landscape leans towards controlling the extrinsic parameters [219, 208]. However, there exists work bridging extrinsic and intrinsic parametric control to satisfy the cinematographer. For example, Pueyo et al [195] present CineMPC which searches for an optimum trajectory (for a drone) and camera angle using a model predicted control (MPC) framework. MPC [33] achieves a process output such as a camera action by considering future time instances/horizons and minimising the cost of selecting different actions. CineMPC models a finite horizon which is continually displaced until all actions cease and constrains the objective function to be flexible to different camera configurations and visual aesthetics. The authors curate mathematical expressions to account for composition, depth of field and canonical shots. Thus, the subsequent cost functions can be considered naive approximations of canonical cinematographic style.
While there exists a lot of work on camera calibration for automated control, the landscape with regards to 3-D reconstruction is more fertile. Discussed by Remondino and Fraser [208], the general approaches to calibration are Tsai [251], Heikkila and Silvén [95] and Zhang [294] which model distortion intrinsic parameters for a pinhole camera. Nevertheless, there is no universal or flexible automation scheme that achieves this [208]. This is why we still see splintered research on subsets paradigms such as Pitombeira and Mitishita [191] which focuses on calibrating the zoom and focus features for scenarios where modelling fixed focus and zoom lenses are infeasible. Or Jasińska et al [113], which looks at improving geometric stability of videos using SfM of multi-view stereo (MVS) images for photogrammetric reconstruction.
Interestingly, there is work relating to INR and NeRF modelling which addresses the whole problem [300]. The classical model in Mildenhall et al [166] naively uses COLMAP [226] whereas recent work has focused on joint-optimization [279, 274, 267, 250]. While many of these papers improve upon pose estimation for a fixed pinhole camera model, there is work on evolving the lens model as well. For example, Xian et al [275] presents a formulation of a ResNet-based lens distortion model and robust patern-based calibration to provide a thin-lens model equitable for NeRF reconstruction as well as other vision based tasks.
3.2.3 Neural 3-D Representations of Dynamic Scenes
Reconstructing real scenes as 3-D representations presents a host of new solutions to existing cinematographic problems, such as novel view synthesis for content acquisition, and elevates prior 2-D based paradigms to 3-D. For instance, in Section 3.1 we discuss how 3-D reconstruction could be used for evaluating IC research on DA by providing realistic virtual environments, for testing through simulation.
There are two general formulations of this problem:
  1. 1.
    Monocular scenes captured with a single camera moving around a moving object [196, 278, 182]
  2. 2.
    MVS scenes containing a single action captured using multiple cameras which are often static [217, 31, 141]
There is a third formulation that has been considered [78], forward facing scenes where a single camera is bounded to a single plane of motion. However this has been weakly adopted as a universal paradigm. For example, K-Planes [78] mimics the general consensus on using normalized device coordinates (NDC) and scene-contraction used in Barron et al [16] to “hack” at this problem. However, this offers no ability to render 6 degrees of freedom (6DoF) dynamic video. Doing so would require hallucinating obstructed geometry (the ‘ “behind” of a scene) which introduces a whole new paradigm.
Methods that attend to the general formulations are shown in Table 2. These methods are not the only concurrent solutions, nonetheless they represent the wide variety of available solutions.
Method
Speed
Commercial GPU
Explicit
Ancestors
D-NeRF D NeRF
Dy-NeRF D NeRF
V4D h -
NeRFPlayer h Instant-NGP, TensorRF
K-Planes m NeRF-W, Instant-NGP, DVGO
HyperReel h TensorRF
NeRF-DS h HyperNeRF
HexPlane h TensorRF
Tensor4D h NeRF-T, D-NeRF
DynIBaR D NSFF, IBRNet
Table 2: Overview of Dynamic NeRFs: Speed measures the training time where D is days, h is hours and m is minutes. Commercial GPU indicates whether the authors tested with commercially available GPUs. Explicit indicates where an explicit representation was used to boost training and inference speed. Ancestors indicates the prior work influencing each method.
There are two traits that currently differentiate originality of research. The first is the proposal of new space-time representations. The second is the manipulation of representations to enhance the learning of temporal elements.
One of the earliest and easiest-to-understand proposals for a new space-time representation is D-NeRF [196]. This method models the deformation field $\Phi(\textbf{x},t)\rightarrow\Delta\textbf{x}$ where $\Delta\textbf{x}$ is the predicted positional change of a ray-sample relative to a canonical static field $\Phi(\textbf{x}+\Delta\textbf{x},t)\rightarrow(\textbf{c},\sigma)$. Relative to global space, this learns an SE(3) transformation. Learning a canonical static space is a robust way of ensuring volumetric consistency with time. Though intuitive, this lends itself to issues when a scene is not continuously in-frame. To counterbalance this, methods such as K-Planes decompose static and dynamic volumes representations. More specifically, K-Planes does this by projecting ray samples (containing $(\textbf{x},t)$) into 6 feature planes, three representing static space and three representing dynamic space. The inputs are normalized between $[0,N]$, projected onto the feature planes and bi-linearly interpolated among varying scales (i.e. coarse and fine features). To decode the features, attaining $(c,\sigma)$, element-wise multiplication is used to recover a final feature vector which is passed into a feature decoder (for explicit representation) or small MLP (for implicit representation). Overcoming similar problems with large and slow-to-train dynamic representation we see GS alternatives such as 4DGS [271] [^12] Pre-print available on arXiv, which use the same K-Plane decompisition, though use it to derive covariance and visual properties, rather than visual-only properties as with K-Planes.
Another interesting solution to the issue of volumetric consistency is key-frame interpolation of static fields. HyperReel exemplifies this by learning the displacement vectors and geometric primitives of a jointly learned key-framed static field. To learn many key-frame fields for a single video, HyperReel builds upon TensorRF due to its compact nature and fast learning ability. Furthermore, as this only provides a discrete set of time-dependant snapshots the authors propose modelling the velocity of volumes at each key-frame. Similarly to the principles of D-NeRF, this enhances the temporal quality of the radiance field representation.
Overall, methods are only modestly capable of modelling dynamic scenes with 6-DoF and usually require additional modification to handle non-generic scenes, such as forward-facing [10, 78]. Despite this, we still find work that tends to production specific needs such as editable NeRFs and GSs’ representations [136, 298, 109]. For cinematographers this means patience as we continue to witness increasing interest in this field.
3.2.4 Neural 3-D Representations of Humans
Modelling non-rigid or deformable geometry, for example a human, is a classical problem for graphics research. This topic is well aligned with research on dynamic NeRFs and broadens human-centered computer vision research to the 3-D case.
Similarly to the prior subsection, the two general problem formulations are multi-view and monocular scenes, which we use to distinguish current research objectives. We can additionally differentiate work by it’s reliance on generalising humans and their poses. HumanNeRF [296] exemplifies both of these points as a method that focuses on the multi-view paradigm while generating generalizable human poses. The authors propose using a NeRF to learn the skinner multi-person linear model (SMPL) [153], i.e. the geometry and appearance of an actor. These features are learnt and used to train a novel neural appearance blending field. Similarly to D-NeRF, the generalizable NeRF learns the canonical and deformation field of an actor by taking inputs of an SMPL skeleton and pixel aligned features and outputting the colour and density of a sample. The appearance blending field refines texture details by accounting for the colour of aligned features from neighbouring views. This model performs well, however it requires carefully placed cameras and struggles with new poses.
MonoHuman [286] overcomes this by using a shared bidirectional deformation modules that disentangles forward and backward deformation into rigid skeletal motion and non-rigid motion. Forward deformation regards the transformation from canonical space to a unique observation space, while backward deformation accomplishes the opposite. To guide training for new poses, forward correspondence features at known key frames are selected from an observation bank and visual features are evaluated relative to the features in the new observation space. This improves volumetric consistency between different observation spaces, meaning sequences of actions can be recovered with more confidence. Additionally, the method is less vulnerable to issues with using monocular video as the observation bank can be used to improve reconstruction of new non-rigid actions.
In cinematographic practice, neural human modelling methods relives us of the accessibility constraints of motion capture (MoCap) suits [260, 261] - which are currently viewed as the gold standard MoCap systems. The authors of the neural motion (NeMo) model [261] exemplify this by testing their proposed framework on athletic actions, taken from the Penn Action Dataset [292]. Like MonoHuman and HumanNeRF, NeMo generalizes motions using multi-view video. Though notably this is achieved by inputting videos of the same action with varying scene conditions, such as different actors and lighting. To handle un-synchronized actions a phase network is introduced as time-based warping function to align poses (i.e. joint angles and translation of motion) in an action sequence. Furthermore, scene-specific feature embeddings handle the visual differences between scenes. Consequently, the phase and instance features are used as inputs to the NeMo network where the outputs define the joint angles and translation of motion - used to render an SMPL model.
Conclusively, this field of research is popular and highly relevant to IC. While it is currently limited by training and inference time [261, 286] there is potential for it use in applications like MoCap, which may be additionally beneficial for actor-based ICVFX.
3.2.5 Challenges and Future Work
With ICVFX, research is limited by access to expensive technologies such as cameras and LED volumes. Considering the novelty of this production format, we are certain to see more research emerge as more stages are built.
Aside from current research on improving the quality of stage lighting, there are other areas that we believe are worth pursuing. LED volume productions (or virtual stage production (VSP)) is heavy on pre-production planning, thus there is potential for supporting tasks like pre-production shot visualization, or 3-D stage design. This area of research branches well with 3-D capture and modelling research. Real-time 3-D view-synthesis (rendering) could also play a role in in-camera background projection for VSP. Though prior to this there are steps yet to be taken in providing high-resolution renders or real-time scene changes.
With the recent introduction of open-source pipelines for NeRF [243, 140] modifying and testing different NeRFs is trivial. However, there is a lack of understanding when it comes to the optimal strategy of capturing media for training NeRFs. Additionally, current methods for evaluating NeRFs using 2-D image-based metrics such as PSNR, SSIM and LPIPS are perhaps not completely reflective of the accuracy of the 3-D fields which we expect a NeRF network to learn [243, 83]. This is specifically true for the dynamic use-case. Therefore, we believe this field still has a considerable journey ahead before we can confidently select models for large scale commercial productions. It is also worth noting that automated 3-D capture technologies play a great role in the general worry over the ownership of an actors visual essence. In Section 4.1 we elaborate on the social responsibility of researchers as we find this ethical dilemma akin to issues previously faced with Deep Fake research.
3.3 Live Production
IC research for live production is conditional on the type of environment, types of actions expected to be displayed (e.g. routines or phases in a performance or sport), location of audience members and the focal point for the digital audience. Analysing the current play or performance can live selection of cameras and shots. Additionally, challenging environments and scenarios may lead to degraded visual quality, thus real time solutions are required for image correction. For example, image analysis from a game of football is challenged by noise arising from spectators or advertisement banners [187, 238, 270], while water-sports will face problems of distortion from partially underwater participants, light artifacts from surface reflection and noise from turbulent water [270, 104]. In this section we link work concerning more research on object tracking and human pose estimation and specifically highlight cases that could inform live shot selection or real-time image correction methods.
3.3.1 Human Pose Estimation
Human pose estimation (HPE) is a popular paradigm in computer vision research [11, 223, 297]. We see cases of HPE used in cinematic scene analysis [265, 273] and also in live event broadcasting. For example, for sport events analysing a player’s pose may allow us to to forecast a series of entertaining events which could provoke particular cinematographic shots. For sport there is ultimately one major factor that divides research: whether or not general application is possible. This is an essential delineation to make prior to reviewing a methods [11] as solutions are heterogeneous [^13] E.g. anatomical poses will vary between sport.. Aside from this there are special cases where models take-on specific challenges, such as exploring HPE in 3D [265, 237], in a highly dynamic environments [98] or for team sports [27].
Hu [105] looks at real-time football player (single-target) posture estimation for live presentational analysis. To accomplish this, confidence weighted fusion and visual attention is used handle problems with colour camouflage and static foreground features, to first identify the target foreground features. Figure13 illustrates pixel-based joint verification for identifying key-point target features, using local binary similarity patterns (lbsp). Then a heat map is generated using a ResNet, a stacked hour glass network and deep-layer aggregation (DLA). The DLA collects features from CNN layers to determine the shared features, i.e. it aggregates features at each layer relating to a verified joint and tries to classify what it is (knee, elbow, etc.).For optimization, the model is updated with adaptive strategies relating to the confidence weight of each pixel and their corresponding weighted fusion sum - based on the joint classification. This methods shows significant improvement in joint and posture detection compared to other state-of-the-art methods [257, 190, 103, 184].
(a) Joint-verification using color features, $v_{t}$, and the local binary similarity pattern, $lbsp_{t}$, of pixel $x$ at frame $t$. $R_{color}$ and $R_{lbsp}$ are predefined distance thresholds and $\{v_{1},v_{2},...,v_{n}\}$ and $\{lbsp_{1},lbsp_{2},...,lbsp_{m}\}$ are a set of known color and binary similarity features. The distance function between the pixel’s colour feature, $lbsp$ feature and features in their respective sets, is evaluated w.r.t the thresholds and the binary result is fused using a logical AND operation to determine where a pixel belongs to a joint.
(b) Local binary pattern example, where a binary similarity function evaluates a given pixel with pixels that fall on it’s local binary pattern. The similarity is taken w.r.t to a threshold, $T$.
Figure 13: Per-pixel binary joint verification used for identifying pixels for joint classification.
Differently, Bridgeman et al [27] looks at tracking and analysing multiple targets simultaneously. This method builds upon existing research by Cao et al [36]. The goal is to construct a 3D (skeletal) pose model using multiple 2D views, while tackling prior issues of long processing time and reliance on appearance-based queues for initiating feature detection. This is achieved by presenting a greedy approach to multi-view detection. The approach has three steps: (1) 2D Pose Error Correction is done by flipping incorrectly oriented body parts and dividing by body part, (2) Per-frame 2D pose association determines a consistent label for each body part across multiple views, found greedily by selecting the best pose from a weighted rank, and (3) 3D skeleton tracking uses the 2D labelled poses to generate a 3D skeleton for each individual. The novelty in this paper is it’s use of multi-view HPE to verify joint placements in 3D.
With similar ambitions, Song and Fan [237] explicitly improves pose detection methods by using different modules from the VGG11 network for different feature fusion methods. To accomplish this, estimates of features are sampled from the VGG11 network and local features of points are first aggregated. After the image passes through a semantic segmentation network, a segmented target (i.e. a body part) is passed into the feature fusion network. This splices and fuses the RGB features of the segmented image, local point-cloud features and global view features to form a final feature vector which is then reduced to a scalar value for classifying the body part.
Wang et al [258] looks at HPE using fixed aerial cameras. As with previous models, a CNN is used for target (player) detection and a YOLOv2 model is used for pose detection, trained on public aerial image data sets. Processed images undergo further classification of normal and abnormal poses. Conclusively, the model provides an insightful way to detect posture from a common viewing angle for events, though is limited by its ability to detect abnormal poses consequent of a biased data set - it is challenging to find publicly available data sets pertaining to abnormal poses. It would be interesting to see this developed for real-time application, particularly as wireless video transmission protocols are shifting to improve on wireless-latency and error handling [3].
3.3.2 Object Tracking
As with HPE, object detection and tracking (ODT) has a significant presence in sports broadcasting, for similar reasons. Unlike HPE, solutions are not prescribed on a case-by-case basis. Rather they tackle implicit issues given a set of idiomatic environment characteristics [124]. For example for football, we could envision a method for detecting a ball in motion within a noisy environment. This translates well to other sports with similar tropes, such as handball.
[124] presents a lengthy review of ODT approaches including evolutionary algorithms, conventional neural networks and feed-forward neural networks (FNNs). It is shown that there are suitable approaches for almost any combination of known visual constraints. For example, FNNs are most successful for video streams with irregular illumination change and noise, whilst darker hues and low frame rate are best handled by DL-based CNN classifiers.
Within IC we find two distinct paradigms, (1) singular target ODT and (2) multiple target ODT. (1) is much easier to solve and solutions tend to use YOLO detection. For example, Wu et al [270] looks at two sports in particular, swimming and table tennis for over-the-top and in-stadium large-screen broadcasting. With the table tennis case study, problems relate to resolving high-speed motion blur for small moving objects to capture the game live in 3-D. While with swimming, problems relate to light-reflection obscuring the cameras vision of swimmers and water occlusion. For table tennis, a YOLOv4 model is used to estimate ball-bounce from a (singular view) video stream and uses this to define a 3-D trajectory. For swimming the solution is a reduced ResNet-50 neural network model acting as the base-network, alongside a modified SiamRPN++ model for swimmer-tracking. To support recognition the camera’s view is masked to suppress background noise. Wu et al [270] concludes on difficulties with the communication of heterogeneous interfaces between collaborators as well as providing the feasibility of this model in actual practice.
As concluded by Burić et al [32], YOLO methods are the fastest to test and fine-tune [^14] Though, we note only YOLOv2 was tested in this review thus is an attractive case for practical development. They are proven to be sufficient particularly for occlusions. The less successful alternative, Mask R-CNN [93], uses two networks, one for detecting regions of interest (i.e. a Regional Proposal Network (RPN)) and a deep CNN for determining how likely a group of pixels within a region of interest is associated with an object.
On the other hand, point (2) concerns multiple target ODT which we split into two further sub-objectives: (i) ODT for objects of the same family, and (ii) ODT for objects from different families (such as player and ball detection in Moon et al [169]). Şah and Direkoğlu [218] reviews several popular methods for in-field-sports broadcasting. Conventional approaches detect visual changes using classical image processing methods [69, 13]. Though, these solutions are naive to several constraints such as (ii). Alternatively, DL-based methods are usually constructed as Fast R-CNNs [86] (ancestor of the Masked R-CNN) and YOLO [193, 203, 210, 111]. Yet again, there is also a case to be made for Masked R-CNN [293]. Surprisingly, Şah and Direkoğlu [218] finds that conventional methods outrank DL approaches - though not by much. It is reasoned that with a larger number of targets to track, deep CNNs are unable to distinguish lower resolution features whilst conventional methods are totally reliant on high-resolution imagery. In the case of (ii), DL approaches are currently unfeasible as small objects detection presents further issues tied to low resolution feature detection [218].
3.3.3 Challenges and Future Work
Relative to other forms of production, live broadcasting has a long standing relationship with AI research. We see this through considerable efforts to elevate the efficacy of pose and object detectors for human targets. We have doubts on the current practicality of these models considering their lack of accuracy relative to human ability, though this is well understood by research and often the main objective. Therefore, we do not expect future work will deviate from investigating this paradigm. Considering the number of processes involved, there are several other feasible avenues for research. For example, we could extend the supporting infrastructure by investigating different masking strategies for suppressing the general set of noisy surroundings. We could also investigate image detection and correction methods of visual artifacts, such as the reflection of light from a swimming pool [270].
Furthermore, a significant sum of footage is readily accessible online. Granted, there are biases in the popularity, gender and racial diversity of sporting events, which should be avoided where possible. Nonetheless, this indicates a need to source more diverse data-sets for general use.
Otherwise, we are interested to see how live IC could play a larger part in driving content capture. For example, we are interested to see more research in optimising shot composition for anticipating events. It would also be exciting to see stylized shot compositions that look at capturing in-game events with a variation of techniques. For instance, using cinematic style transfer [265] to replicate praiseworthy shots. Perhaps one day we may be able to watch a sports game in a Tarantio-esque style.
3.4 Aerial Production
Similar to many other applications of IC, the problem landscape for UAV-based cinematography is generally undefined [156]. This is further pronounced when we consider the limitations that come with commercial drones, which are not easy to outfit for cinematographic purpose. Mademlis et al [156] produces an interesting review of concurrent challenges, outlining that due to badly defined problem space, lack of technological accessibility as well as roughly drawn legal, ethical and safety constraints, it can be challenging to find a solution which fits the general needs of cinematographers. Despite this, work on autonomous UAVs continues to grow and more recently we find large variety of seemingly successful approaches to UAV control for IC.
In this section we make the key delineation between single drone and multi-drone systems. This is important to consider for cinematographers with tight budgets but also as the technology between the two can greatly differ.
3.4.1 Automated Single UAV control
Aside from budget constraints, the choice to use a single UAV may come down to the size and accessibility of the camera’s viewing search space. For example a small viewing region around an actor to capture close to mid range shots may only require a single versatile drone. With limited airspace around a subject more drones demands additional constraints to avoid crashes further diminishing the possibility of real-time solutions. Thus there is an area of research on single UAV control referred to as autonomous “follow me” quadrotors [118]. There are a number of vision and sensor-oriented solutions within this domain [173, 246, 145, 60] as well as drone manufacturers that supply this feature [118], like the 3DR Solo and DJI Phantom.
To accomplish this work there are a number of physical and technical limitations to consider. The first is the trade off between online and offline solutions [26, 197]. Online solutions [107, 26] make decisions quickly, responsively and are pragmatic when active elements in a scene move in unpredictable ways, such as an actors movement. Additionally, work exists on physical camera and drone drone modification for improved online and onboard capability [107]. Offline solutions can enable solving more complex challenges, such as swarm robotics [158].
The second challenge to consider is the method of detecting actors. This is useful for trajectory planning but also gimbal control. Bonatti et al [26] considers using a single shot detector [151], such as YOLO9000 [206] and Fast R-CNN [210], for a problem demanding smooth gimbal control and determines that Fast R-CNN is optimal given the speed-accuracy trade off for defining a trajectory. They additionally use MobileNet to perform low-memory feature extraction with fast inference speed. Finally, for indefinite actor tracking they use a KCF [99], taking advantage of the fact that some learning algorithms perform better in the Fourier domain. The KCF tracker relies on kernel ridge regression which is a kernalized version of the linear correlation filter, forming the basis for the fastest trackers available [24, 23].
Finally the third and most evident challenge is how trajectory planning is handled with regard to certain cinematographic objectives. A tempting option is to adapt cinematographic concepts to mathematical expressions that are subsequently optimized for control [195, 118]. Alternatively, Ashtari et al [9] proposes a method that models the dynamic motion of a human camera operator in real time. The authors follow works on modelling the vertical and lateral displacement of a walking patterns [39, 302] combining approaches into a single routine that additionally considers the rotation of the drone and damping effects to simulate different types of camera equipment. Another approach is to let a reinforcement learning agent control the camera motion. Gschwindt et al [88] builds on CHOMP [204] to parameterize smooth trajectory planning while using a deep Q network (DQN) to lead target shot selection. Two methods are proposed for training the DQN. The first method is a human-crafted reward function; like adapting cinematographic shots to optimization functions, this reward function accounts for the actors presences in a shot, the duration and the shot angle. The second method is human-led observation which rewards the DQN relative to the cinematographers subjective opinion.
Ultimately, the design of autonomous single-camera UAV systems is ambiguous. For cinematographers, this means there are a number of trade offs to consider, such as the complexity and feasibility of using more flexibly systems (e.g. drones with gimbals or ”follow me” modes). We agree with the premise set by Mademlis et al [156], that work still needs to be done unifying the objectives of autonomous drone systems. However we also believe that a universal paradigm could inhibit varying uses, for example mimicking human walk patterns compared and optimising shot composition present different objectives and outcomes.
3.4.2 Automated Multiple UAV control
For multi-UAV technologies, the limitations of single UAV control are amplified. For example, communication, cooperation, finite bandwidth and safety concerns become more precarious to manage as the number of UAVs increases [157]. This is one reason behind the use of schedulers [37, 248]. Scheduling drones can insinuate two types of tasks, a swarm of cooperating drones with a unified purpose and non-cooperative drones [158]. For cinematographers, this is comparative to a drone fleet being used to capture multiple views achieving the same cinematographic objective (for example, a thematically driven system [172]) or a fleet used to capture a varying selection of shots for sport or event like filming, where a human or autonomous director makes the final decision over which shot to broadcast [37].
Additionally delineating from single UAV capture, it becomes less plausible to use GPS-based localization for trajectory planning [157]. This is because GPS systems introduce noise during drone localization which perturbs drone formations that demand accurate localization. Furthermore, a similar problem is faced with SLAM-only methods for localization. Hence, an inertial measurement unit (IMU)-GPS-SLAM fused system [157] has been employed. This problem is also shared with works that lie outside of the IC scope [277, 90, 66]. For example, Yan et al [277] looks at a fusing a VO [^15] Discussed in Section 2 system with IMU integrated on an extended Kalman filter; inspired by Bloesch et al [21]. With the additional curation of a dense 3-D map of the target environment, built using ORB-SLAM [263], automatic robotic localization and control in an agricultural context (i.e. inside a green house and outside in a field) is made possible. Interestingly, the authors experiment on cluttered and non-cluttered environments which is a shared characteristic of production environments - namely the variation of clutter between different sets.
Nägeli et al [172] directly address this problem and further highlight challenges with dynamic entities (drones, humans, moving set etc.), using model predicted contour control (MPCC) [^16] This is related to MPC, previously mentioned.. Confronted with a highly constrained scenario, the authors simplify the problem using manually defined “virtual rails” for each drone. This acts as a coarse trajectory guide that drones loosely follow to avoid collisions while achieving one-shot aesthetic objectives. To track and update the state of moving targets a Kalman filter is used. Furthermore, the authors employ an actor-driven view-framing method which can be adjusted in real-time via general user interface (GUI) by a cinematographer for varied framing. While not explored in the paper, this could lead to extensions that look at varying the transition between framing inputs at different times to produce aesthetically pleasing transitions. Finally, overcoming the collision constraints, the authors model regions to be avoided as ellipses around a subject as a hard constraint (i.e. high penalization), using slack-variables to indicate when a horizon is foreseeable. Hence, when slack variables are high, the problem is deemed infeasible either due to violating the collision constraints or the computational budget being exhausted. We note that slack variables are commonly used in MPC research to model pareto-optimal solutions.
Overall, the limitations raised by researchers are wide-spread. Since each production scenario is unique and solutions must be robust due to safety concerns, there is no work that currently solves the general paradigm. We share our formalization of the general paradigm with Mademlis et al [156], whereby the ideal is a system that can act on a set of high-level inputs from an operator, who is not required to be technically knowledgeable.
3.4.3 Challenges and Future Work
Respecting the goal of a general automated UAV solution for IC, the prior subsections lead us to encourage prioritising the discovery of a holistic set solutions. The breadth of use-cases are wide and often unique, however much like the DA problem (discussed in Section3.1.2), solutions heavily rely on the development of related works. For example, solutions to the UAV localization paradigm canonically involve extensions of SLAM and VO research. With recent developments in camera calibration, consequent to the explosion of NeRF research, we are thus likely to see extended use of visual methods for localization and path generation in the future. Considering this avoids the need for GPS-related hardware, it may result in longer flying time and cheaper budgets which are particularly useful for multi-UAV paradigm.
Moreover, with new developments in dynamic and human NeRFs, discussed in Section 3.2.3 and 3.2.4, we see additional potential in training and/or testing automated drones on unforeseen circumstances using real-world digital twins. This could be used to accomplish safety tests or to train UAVs on scene-specific scenarios. For example, if we wanted a UAV to learn to react to actors disobeying drone safety protocol, we could train a NeRF environment using MonoHuman and interpolate sequences of human poses which defy protocol.
On the other hand, the steady developments of NeRF could render UAV-based solutions redundant for directly acquiring the desired content. Rather, it is likely that UAVs become an intermediary step for NeRF-like content acquisition, as it separates the problem into two steps: (1) UAV image acquisition with the objective of optimising a NeRF with 6DoF, and (2) content acquisition within NeRF with 6DoF flexibility. This simplifies the current paradigm, which accomplishes localization and path generation constrained by subjective visual objectives.
Alternatively, while this may apply to the general UAV use-case there are situations that still require attention from IC researchers. For example, the follow-me paradigm implies the use of UAVs for direct content acquisition and relies more on the ability of a drone to move with a visual anchor, like a human target, under cinematographic objectives, like a desired walk pattern or jitter. Considering this is a simpler problem with arguably easier to define objectives, the intervention of NeRF may be unnecessary.
4 General Remarks
This literature review begins with the characterisation of IC as the field of research dedicated to exploring AI solutions for cinematographic video production. We subsequently reviewed four production mediums that we believe are most relevance to current researchers and interested cinematographers, including a subsection on general production tools. This leads us to present our final remarks.
There are certainly several favourable fields of research that have potential to shape the industry going forward:
  1. 1.
    Computational Language Structures, discussed in Section 3.1.1
  2. 2.
    LED Volumes, discussed in Section 3.2.1
  3. 3.
    NeRFs, discussed in Section 3.2.3 and 3.2.4
With regards to computational language structures, we do not believe that the current state of research is telling of its potential impact on IC. Instead, we believe with the current state of research on LLMs and it’s use in the public domain [^17] E.g. generating trailers from text/scripts [211, 54], that natural language decision making for IC will be facilitated. Though we have yet to see this is investigated in research, we believe there is cause to pursue its application in fields like automated directive assistance, MP for planning and automated UAV control.
With regards to LED volumes and NeRFs we believe the impact on industry has already been felt. For the former, the adoption of this new technology has offered radical change to canonical chroma-keying practices and offers a new image-based paradigm for computer vision researchers. Regarding the latter, research has demonstrated a flexible pool of new solutions to problems which were previously solved through with sculpting/3-D modelling tools and/or hardware dependant photogrammetric methods. With the current push for compact, fast and accurate representations, we would not be surprised if production ready NeRFs became available soon.
Conclusively, we hope that our work inspires more reviews on future developments of IC research. The fields of research and the relevant video production industries we have discussed are highly susceptible to change, as exemplified by the influence of the LED volume. Thus we should acknowledge the need for concurrent reviews. Finally, to further bridge the gap between IC researchers and industry professionals we encourage future reviews that targets industry professionals directly.
4.1 Social Responsibility
The impact of evolving technologies on the creative industries is multifaceted. Regarding IC, there are two matters that underpin widespread concerns with freely advancing AI: (1) replacing actors with human-like AI models [17, 51, 52] and (2) using generative AI and/or 3-D models to produce pornographic content [245, 178]. The matters mainly concern fields such as automated 3-D capture and generative modelling, where the ability to capture a persons physical essence is currently possible to a high degree for images and likely soon-to-be for videos. As we discuss these topics in Section 3.2.4 and 2.2, respectively, we take the opportunity now to provide awareness on the issues and discuss steps which can be taken to avoid certain ethical dilemmas.
Replacing real actors can be achieved by either modelling a real human or generating a virtual human [17]. For actors this is problematic as current stars fear abuse of existing content containing detailed depictions of themselves. This may also lead to marginalizing real people for virtual celebrity look-a-likes.
The adult entertainment industry is afflicted by similar problems, though these are enhanced by a comprehensive set of concerns including the widened accessibility of mature content from minors and the treatment of women in the work place [85, 239]. Notably, (mis-)use of deep-fakes leads to serious outcomes such as the fact that 96% of deepfakes are sexually explicit depictions of women who didn’t consent [245] and also has profound implications in-sighting cases of child pornography [178]. Embellishing the issue further, Kirchengast [128] reviews current regulatory mechanisms and legislative powers in the US, concluding that present solutions require thorough critique and serious investigation. This is discussed in consideration to the current level of criminalization, which weakly implicates the parties at fault.
Consequently, we believe researchers should take steps to avoid contributing these problems. The general choices include, highlighting awareness as we have in this section, limiting the use of publicly available code through licensing and applying due-diligence when collaborating with external partners. We also believe that less pressure should be placed on researchers that choose to publish with close-source code. Hence, we should learn from prior instances involving generative AI and avoid publishing state-of-the-art models (and optimised parameter sets) that do not have robust measures to counter misuse. Furthermore, this could be accompanied by statements of transparency, outlining the efforts that have been made to avoid code misuse. Researchers may also consider the chain of fault and potential legal ramifications that comes with irresponsibly publishing code. As legislative discussions have not concluded on a system for assigning fault, this is worth considering.